# A Data-Aware Fourier Neural Operator for Modeling Spatiotemporal Electromagnetic Fields

### v1

@@ Abstract
Neural operators have recently emerged as a powerful tool for solving partial differential equations (PDEs), including Maxwell’s equations in computational electromagnetics. Prior machine learning models often fail to capture both temporal field evolution and generalization to irregular geometries. Here, we introduce a Data-Aware Fourier Neural Operator (DA-FNO) as a surrogate solver. Applied autoregressively, it predicts the temporal evolution of all field components while monitoring energy dynamics, terminating automatically once energy converges. The model generalizes to complex geometries and the optical C-band without retraining, achieving a $7.5\times$ speedup with nearly $92.5\%$ accuracy. This approach offers a potentially efficient and accurate alternative to conventional iterative solvers for C-band photonic simulations.
@@ 1 Introduction
Solving Maxwell’s equations remains a fundamental challenge in photonics. Current engineering practices predominantly rely on numerical methods such as the Finite Element Method (FEM) [1, 2], Finite Difference Frequency Domain (FDFD) [3, 4], and Finite Difference Time Domain (FDTD) [5, 6] algorithms. Although full-wave Maxwell solvers provide accuracy, their substantial computational costs restrict the scalability of applicable scenarios. Among these methods, the FDTD algorithm stands out for its explicit temporal evolution of electromagnetic fields, enabling efficient time-resolved simulations such as ultrafast optics and broadband excitation. However, its strict adherence to the Courant-Friedrichs-Lewy (CFL) stability condition imposes stringent spatiotemporal resolution constraints, leading to excessive computational costs for high-precision or large-scale simulations.
In recent years, machine learning (ML) approaches have been extensively explored for solving electromagnetic field problems, often demonstrating significantly faster computational speeds than traditional methods while maintaining acceptable accuracy. Convolutional neural networks (CNNs) with their local receptive fields and various recurrent networks, such as Recurrent CNN (RCNN) [7, 8, 9] and Long Short-Term Memory - FDTD (LSTM-FDTD) [10, 11], have been widely applied to improve the FDTD algorithm. Recent extensions also incorporate transformer architectures [12, 13, 14] to model long-range electromagnetic interactions through self-attention mechanisms. However, these conventional network architectures struggle to generalize across irregular and arbitrary geometries despite their iterative temporal prediction capabilities. In contrast to these data-driven approaches, physics-informed neural networks (PINNs) [15, 16] have emerged as an unsupervised approach that embeds physical laws into the loss function, enabling PDEs solving without labeled data. Concurrently, neural operator methods have gained attention for learning mappings between function spaces, effectively modeling the underlying PDE operators. Representative models such as Deep operator networks (DeepONets) [17] and Fourier neural operators (FNOs) [18, 19, 20, 21] have been applied to solve Maxwell’s equations, demonstrating promising results in computational electromagnetics. Nevertheless, most operator-based frameworks and PINNs focus exclusively on frequency-domain field predictions, neglecting time-domain dynamics, while the few capable of temporal outputs [18, 19] are confined to parametric variations of regular geometries.
@@
In this work, we propose a Data-Aware Fourier Neural Operator (DA-FNO) that is applied autoregressively as a surrogate solver for 2D electromagnetic field simulations. The proposed model is capable of generalizing across unseen geometries, targeting Transverse Magnetic (TM)-polarized plane wave scattering in a simulation domain. Similar to the FDTD method, the model iteratively outputs all field components $(E_{z},H_{x},H_{y})$ over time, allowing for dynamic tracking of the scattering process. Therefore, the DA-FNO model simultaneously supports the temporal evolution of electromagnetic fields and generalization to complex and randomized geometries. In addition, it demonstrates strong wavelength generalization across the optical communication C-band. The model monitors the instantaneous energy within the domain, progressively reducing energy as the outgoing scattered waves evolve, and automatically terminates the simulation once the energy falls below a convergence criterion, $\delta$, of its historical maximum. Moreover, it eliminates CFL constraints, enabling larger time steps ($\Delta t$), achieving a $7.5\times$ speedup over the conventional FDTD method while maintaining $92.5\%$ solution accuracy when $\delta=10^{-2}$. Fig. 1 illustrates the schematic diagram of the proposed model. The FDTD simulation is performed with a time step of $\Delta t$. The DA-FNO model outputs the three field components ($E_{z},H_{x},H_{y}$) of the TM-polarized wave at every $m\Delta t$ interval, enabling faster simulation compared to the FDTD. This process continues until the energy reaches convergence.
Figure 1: Schematic diagram of the proposed model. The time step of FDTD is $\Delta t$ while that of the DA-FNO model is $m\Delta t$, enabling faster simulation.
@@ 2 Model Construction
$v_{m+1}\left(x\right)=\sigma\left(\mathcal{F}^{-1}\left(R_{m}\odot\mathcal{F}\left(v_{m}\left(x\right)\right)\right)+W_{m}v_{m}\left(x\right)\right),$ (1)
@@
A pictorial representation of our model architecture is presented in Fig. 3(a). The input tensor of the model stacks five consecutive temporal states $\left\{s_{t-4},\ldots,s_{t}\right\}$ along the channel dimension, with each state comprising electromagnetic field components $\left(E_{z},H_{x},H_{y}\right)$ for the TM-polarized wave. Spatial coordinates $\left(x,y\right)$ and the permittivity distribution $\epsilon$ are first appended to the channels to encode geometric priors. Next, the augmented input is lifted into a high-dimensional latent space via a fully connected layer $P$, and subsequently transformed through four Data-Aware Fourier layers. Then, a linear layer $Q$ maps the refined latent features to the output temporal state $s_{t+1}$, representing the electromagnetic field updates. Finally, the oldest state in the input tensor $s_{t-4}$ is discarded, and the predicted state $s_{t+1}$ is appended to form the updated input tensor $\left\{s_{t-3},\ldots,s_{t+1}\right\}$ for the next iteration. This auto-regressive process continues, with the electromagnetic energy of each newly predicted state being computed as:
@@
Figure 3: (a) The full architecture of the DA-FNO model. The five input temporal field states $\left\{s_{t-4},\ldots,s_{t}\right\}$ is augmented by $S$ with permittivity distribution $\epsilon$ and spatial coordinates $(x,y)$, lifted to a higher-dimensional space by network $P$, processed through four DA Fourier layers and mapped to the next state $s_{t+1}$ by network $Q$. The new state and the last four temporal states of the input form a new input $\left\{s_{t-3},\ldots,s_{t+1}\right\}$, enabling the iteration to continue until energy convergence. (b) The architecture of the DA Fourier layer. The convolutional processing establishes correlations among the three field components. Data-Aware mode selection is performed in the spectral space.
@@
However, when directly fed into the vanilla FNO, these physically coupled field components are treated as decoupled channels during spectral processing. As a result, the vanilla FNO fails to capture the intrinsic physical correlations and constraints dictated by the curl equations, which link the evolution of one component to the spatial variations of the others. To address this architectural limitation and enforce physics-compliant interactions, we introduce three learnable $3\times 3$ convolution kernels before the Fourier transform $\mathcal{F}$ in each Fourier layer, as shown in Fig. 3(b). Here, $T$ denotes a tensor shape transformation that maps the dimension of the field component to the dimension of the feature, while $T^{-1}$ represents its inverse. Each convolution kernel takes all three field components as input channels and produces a single output component. This design allows each field component to dynamically incorporate information from the other two through localized spatial mixing. We adopt convolution rather than other linear transformations because it effectively captures local spatial dependencies, closely resembling the computations of curl operators. Combined with the Fourier transform, the DA-FNO enables complementary learning of both local and global spatial interactions.
Figure 4: (a) Ez field distributions at $t=5\Delta t,13\Delta t$ and $21\Delta t$. (b) The mode selection regions of the vanilla FNO and DA-FNO. Average of the spectral images of all input data, normalized and displayed in logarithmic scale.
@@
Furthermore, we modify the spectral mode selection strategy within the Fourier layers of the vanilla FNO. The vanilla FNO architecture adopts a rectangular truncation window, $\left[-k_{max},k_{max}\right]^{2}$, to retain low-frequency spectral modes. While this promotes smooth outputs and suppress high-frequency noise, it may be suboptimal for electromagnetic simulations, where field interactions with complex geometries give rise to increasingly rich high-frequency features over time. As illustrated in Fig. 4(a), the $E_{z}$ field distributions at $t=5\Delta t$$13\Delta t$, and $21\Delta t$ exhibit progressively intricate spatial features, accompanied by the emergence of high-frequency components. To address this spectral adaptability gap, we propose a data-driven spectral mode selection scheme, consisting of three steps: (1) Perform 2D Fourier transforms of spatiotemporal field slices $(E_{z},H_{x},H_{y})$ across the training dataset. (2) Apply min-max normalization to each spectrum to mitigate amplitude decay effects arising from temporal evolution. (3) Average the normalized spectra and select spectral modes in descending order of magnitude until the cumulative sum exceeds a predefined threshold $\theta$ of the total integrated magnitude. The indices of the selected modes are recorded and used to define the data-driven sampling region in each Fourier layer. Fig. 4(b) illustrates the selection regions (white areas) of the vanilla FNO and the DA-FNO at $\theta=90\%$, each selecting a comparable number of spectral modes. The final panel depicts the averaged Fourier spectrum over 2700 samples. Compared to the vanilla FNO, the selection region of the DA-FNO exhibits a better match to the spectral distribution, capturing more high-amplitude spectral modes.

### v2

# Accelerated Time-Domain Simulation of Complex Photonic Structures with a Data-Aware Fourier Neural Operator

@@ Abstract
[…]
keywords:
Photonic Simulation , Data-Aware Fourier Neural Operator , Finite-Difference Time-Domain , Light Scattering
[…]
[1]organization=State Key Laboratory of Photonics and Communications, School of Electronics, Peking University, city=Beijing, postcode=100871, country=China
[…]
[2]organization=School of Computer & Communication Engineering, University of Science and Technology Beijing, city=Beijing, postcode=100083, country=China
[…]
[3]organization=Peng Cheng Laboratory, city=Shenzhen, postcode=518055, country=China
@@ 1 Introduction
Solving Maxwell’s equations remains a fundamental challenge in photonics. Current engineering practices predominantly rely on numerical methods such as the Finite Element Method (FEM) [33, 10], Finite Difference Frequency Domain (FDFD) [40, 8], and Finite Difference Time Domain (FDTD) [30, 29] algorithms. Although full-wave Maxwell solvers provide accuracy, their substantial computational costs restrict the scalability of applicable scenarios. Among these methods, the FDTD algorithm stands out for its explicit temporal evolution of electromagnetic fields, enabling efficient time-resolved simulations such as ultrafast optics and broadband excitation [26, 20, 21]. However, its strict adherence to the Courant-Friedrichs-Lewy (CFL) stability condition imposes stringent spatiotemporal resolution constraints, leading to excessive computational costs for high-precision or large-scale simulations.
In recent years, machine learning (ML) approaches have been explored to overcome these computational bottlenecks. To capture the temporal evolution of fields, early efforts utilized recurrent architectures like LSTMs[7] and RNNs combined with convolutions[12]. While successful in modeling time-resolved dynamics, these approaches were often limited in their ability to generalize beyond the simple, regular geometries on which they were trained[36, 23, 39, 6]. Other architectures like Transformers[32, 4, 22, 17] and Graph Neural Networks (GNNs) have also been applied. For instance, GNNs have been used to perform a single-step FDTD update, but error accumulation prevented their extension to full iterative time stepping for long-term simulations[13].
Figure 1: Schematic diagram of the proposed model. Similar to the FDTD method, the DA-FNO model iteratively outputs all field components ($E_{z}$, $H_{x}$, $H_{y}$) over time until the energy converges. By eliminating CFL constraints, the DA-FNO allows for larger time steps ($m\Delta t$), thereby enabling faster simulations.
Concurrently, two other paradigms have gained traction. Physics-Informed Neural Networks (PINNs)[27, 3] offer an unsupervised approach by embedding physical laws directly into the loss function, enabling PDE solving without labeled data. This has been applied to predict time-domain electromagnetic fields[18, 25] and optimize nanophotonic structures like microlenses[16, 5] and metagratings[2]. Separately, neural operators[14], which learn mappings between function spaces, have shown great promise. Models like Fourier Neural Operators (FNOs)[15] and Deep Operator Networks (DeepONets)[19] have been successfully applied to frequency-domain problems[38, 9], such as modeling scattering from irregular structures[1]. A recent FNO-based model[41] effectively models time-domain electromagnetic fields over a fixed temporal horizon, producing predictions in a single forward pass.
Despite these advances, a critical gap remains. Most existing models are tailored to specific scenarios, and while some can handle irregular structures in the frequency domain, there is still no model that simultaneously achieves long-term time-domain simulation while generalizing to complex and irregular geometries.
@@
In this work, we propose a Data-Aware Fourier Neural Operator (DA-FNO) to address this challenge, applied autoregressively to predict the evolution of 2D electromagnetic fields. The proposed model is capable of generalizing across unseen geometries, targeting Transverse Magnetic (TM)-polarized plane wave scattering in a simulation domain. Similar to the FDTD method, the model iteratively outputs all field components $(E_{z},H_{x},H_{y})$ over time, allowing for dynamic tracking of the scattering process. Therefore, the DA-FNO model simultaneously supports the time-domain evolution of electromagnetic fields and generalization to complex and randomized geometries. In addition, it demonstrates strong wavelength generalization across the optical C-band. To further assess the model’s scalability, we also conduct a preliminary investigation on 3D extruded geometries, verifying the model’s capacity to learn volumetric features. The DA-FNO model monitors the instantaneous energy within the domain, progressively reducing energy as the outgoing scattered waves evolve, and automatically terminates the simulation once the energy falls below a convergence factor, $\delta$, of its historical maximum. Moreover, the elimination of CFL constraints enables the use of a coarse time-step ($m\Delta t$), creating a trade-off between computational speed and solution accuracy: increasing the step size (m) improves speed but decreases accuracy. In a representative configuration with $m=15$ and $\delta=10^{-2}$, the model attains an $11\times$ speedup while maintaining approximately 95% accuracy. Fig. 1 illustrates the schematic diagram of the proposed model. The FDTD simulation is performed with a time step of $\Delta t$. The DA-FNO model outputs the three field components ($E_{z},H_{x},H_{y}$) of the TM-polarized wave at every $m\Delta t$ interval, enabling faster simulation compared to the FDTD. This process continues until the energy reaches convergence.
Full access to the implementation code and the first 200 training samples is provided.
@@ 2 Model Construction
$\small v_{m+1}\left(x\right)=\sigma\left(\mathcal{F}^{-1}\left(R_{m}\odot\mathcal{F}\left(v_{m}\left(x\right)\right)\right)+W_{m}v_{m}\left(x\right)\right),$ (1)
@@
A pictorial representation of our model architecture is presented in Fig. 3(a). The input tensor of the model stacks five consecutive time-domain states $\left\{s_{t-4},\ldots,s_{t}\right\}$ along the channel dimension, with each state comprising electromagnetic field components $\left(E_{z},H_{x},H_{y}\right)$ for the TM-polarized wave. Spatial coordinates $\left(x,y\right)$ and the permittivity distribution $\epsilon$ are first appended to the channels to encode geometric priors. Next, the augmented input is lifted into a high-dimensional latent space via a fully connected layer $P$, and subsequently transformed through four Data-Aware Fourier layers. Then, a linear layer $Q$ maps the refined latent features to the output time-domain state $s_{t+1}$, representing the electromagnetic field updates. Finally, the oldest state in the input tensor $s_{t-4}$ is discarded, and the predicted state $s_{t+1}$ is appended to form the updated input tensor $\left\{s_{t-3},\ldots,s_{t+1}\right\}$ for the next iteration. This auto-regressive process continues, with the electromagnetic energy of each newly predicted state being computed as
@@
Figure 3: (a) The full architecture of the DA-FNO model. The five input time-domain field states $\left\{s_{t-4},\ldots,s_{t}\right\}$ is augmented by $S$ with permittivity distribution $\epsilon$ and spatial coordinates $(x,y)$, lifted to a higher-dimensional space by network $P$, processed through four DA Fourier layers and mapped to the next state $s_{t+1}$ by network $Q$. The new state and the last four time-domain states of the input form a new input $\left\{s_{t-3},\ldots,s_{t+1}\right\}$, enabling the iteration to continue until the energy of the new state falls below a convergence factor, $\delta$, of the historical maximum. (b) The architecture of the DA Fourier layer. The convolutional processing establishes correlations among the three field components. Data-Aware mode selection is performed in the spectral space.
Figure 4: (a) Ez field distributions at $t=5\Delta t,13\Delta t$ and $21\Delta t$. (b) The mode selection regions (white areas) of the vanilla FNO and DA-FNO. Average spectral image of all training data, normalized and displayed in logarithmic scale.
@@
However, when directly fed into the vanilla FNO, these physically coupled field components are treated as decoupled channels during spectral processing. As a result, the vanilla FNO fails to capture the intrinsic physical correlations and constraints dictated by the curl equations, which link the evolution of one component to the spatial variations of the others. To address this architectural limitation and enforce physics-compliant interactions, we introduce a learnable $3\times 3$ convolution layer with 3 input channels and 3 output channels before the Fourier transform $\mathcal{F}$ in each Fourier layer, as shown in Fig. 3(b). Here, $T$ denotes a tensor shape transformation that maps the dimension of the field component to the dimension of the feature, while $T^{-1}$ represents its inverse. The convolution takes all three field components as input channels and generates three output components correspondingly. This design allows each field component to dynamically incorporate information from the other two through localized spatial mixing. We adopt convolution rather than other linear transformations because it effectively captures local spatial dependencies, closely resembling the computations of curl operators. Combined with the Fourier transform, the DA-FNO enables complementary learning of both local and global spatial interactions, where local interactions reflect fine-scale updates, and global interactions describe the overall long-term field evolution, according to the FDTD update equations[30] (see Supplementary material, Section 1 for more details).
@@
Furthermore, we modify the spectral mode selection strategy to better capture the underlying physics of the scattering process. The vanilla FNO architecture adopts a rectangular truncation window, $\left[-k_{max},k_{max}\right]^{2}$, to retain low-frequency spectral modes. While effective for suppressing noise, this fixed, low-pass filtering is suboptimal for complex electromagnetic simulations. As a wave interacts with sub-wavelength features in a complex geometry, physical phenomena such as light scattering and diffraction generate fine spatial features in the field distribution. In the spectral domain, accurately representing these features requires retaining the corresponding high-frequency components (large k-vectors). As illustrated in Fig. 4(a), the $E_{z}$ field distributions at $t=5\Delta t$, $13\Delta t$, and $21\Delta t$ exhibit progressively intricate spatial patterns, confirming the emergence of these critical high-frequency components over time. To address this, we propose a physically-motivated, data-aware spectral mode selection scheme. Instead of a fixed rectangular window, our method learns the most significant spectral modes directly from the training data by performing three steps: (1) Perform 2D Fourier transforms of spatiotemporal field slices $(E_{z},H_{x},H_{y})$ across the training dataset. (2) Apply min-max normalization to each spectrum to mitigate amplitude decay effects arising from time-domain evolution. (3) Average the normalized spectra and select spectral modes in descending order of magnitude until the cumulative sum exceeds a predefined threshold $\theta$ of the total integrated magnitude. The indices of the selected modes are recorded and used to define the sampling region in each Fourier layer. As shown in Fig. 4(b), this selection region (white area) for the DA-FNO ($\theta=0.9$) deviates significantly from the rectangular window, showing a much better match to the true spectral distribution of the fields. This allows it to capture more high-amplitude spectral modes crucial for accurately resolving the scattering effects.
Table 1: ARL1E of the Six Models ($mean\pm std$)
Model Number of selected modes Train Test
DA-FNO ($\theta=0.9$) 1817 0.108 $\pm$ 0.017 0.235 $\pm$ 0.027
DA-FNO ($\theta=0.8$) 933 0.124 $\pm$ 0.015 0.244 $\pm$ 0.030
DA-FNO ($\theta=0.7$) 481 0.128 $\pm$ 0.006 0.249 $\pm$ 0.018
DA-FNO ($\theta=0.6$) 231 0.134 $\pm$ 0.007 0.253 $\pm$ 0.030
Conv 1922 0.116 $\pm$ 0.011 0.246 $\pm$ 0.040
vanilla 1922 0.221 $\pm$ 0.021 0.970 $\pm$ 0.057